A web Music Store offers two versions of a popular song. The size of the standard version is 2.7 megabytes (MB). The size of the high-quality version is 4.3 MB. Yesterday, the high-quality version was downloaded four times as often as the standard version. The total size downloaded for the two versions was 5373 MB. How many downloads of the standard version were there.
Answer by ikleyn(52781)

.
A web Music Store offers two versions of a popular song.
The size of the standard version is 2.7 megabytes (MB).
The size of the high-quality version is 4.3 MB.
Yesterday, the high-quality version was downloaded four times as often as the standard version.
The total size downloaded for the two versions was 5373 MB.
How many downloads of the standard version were there.
~~~~~~~~~~~~~~~~~~~~

Let x be the number of standard versions downloaded 
      (the unknown number under the problem's question).

Then the number of high-quality versions downloaded is 4x.


Write the total memory size equation

    2.7*x + 4.3*(4x) = 5373 MB.


Simplify and find x

    (2.7 + 4*4.3)x = 5373

         19.9x     = 5373

             x     = 5373/19.9 = 270.


ANSWER.  270 standard versions were downloaded.

Solved.